Use factoring to solve the equation.
7 z^ 2 = 161 z - 924

Found 2 solutions by checkley71, stanbon:
Answer by checkley71(8403)

7Z^2=161Z-924
7Z^2-161Z+924=0
7(Z^2-23Z+132)=0
7(Z-11)(Z-12)=0
Z-11=0
Z=11 ANSWER.
Z-12=0
Z=12 ANSWER.

Answer by stanbon(75887)

Use factoring to solve the equation.
7 z^ 2 = 161 z - 924
----------------
7z^2 -161z +924 = 0
(z-12)(7z-77)=0
z = 12 or z = 11
===================
Cheers,
Stan H.